if 21-2x=-9x then x=
Answer by stanbon(75887)

A rectangular pool is surrounded by a walk 4 meters wide. The pool is 6 meters longer than its width. If the total area is 576 square meters more than the area of the pool, find the dimensions of the pool.
How do they get the answer? I've tried by making Total Area = x^2+6x+576, but have been unable to obtain the answer.
------------------
Pool width: "x" meters
Pool length: "x+6" meters
Area of the pool: x^2+6x sq. meters
-----------------------------
Total width : x+8
Total length: x+6+8 = x+14
Total Area: (x+8)(x+14) = x^2+22x+112 sq meters
------------------------------
Equation:
Total area - Pool area = 576 sq. meters
------------------
x^2+22x+112 - x^2+6x = 576
16x +112= 576
16x = 464
x = 29 meters (pool width)
x+6 = 35 meters (pool length)
-------------------------------
Cheers,
Stan H.